OpenStudy (anonymous):

x-c/2 = d solve for x

OpenStudy (anonymous):

\[\frac{ x-c }{ 2 }=d \] solve for x

OpenStudy (anonymous):

All you have to do is to add c/2 to both sides, since that's the inverse of subtracting: x = d + c/2

OpenStudy (anonymous):

you just do the opposite of each

OpenStudy (anonymous):

\[x=d+\frac{ c }{ 2 }\] is that the answer joker2.0